OpenStudy (kj4uts):

A sample of 64 was taken from a population, with the mean of the sample being 119 and the standard deviation of the sample being 16. What is the 68% confidence interval for the mean of the population? A. (115, 123) B. (113, 125) C. (117, 121) D. (118, 120)

OpenStudy (rational):

68% is within 1 sigma of mean, so simply find the standard error and work the confidence interval

OpenStudy (rational):

standard error = 16/sqrt(64) = 16/8 = 2

OpenStudy (rational):

therefore the confidence interval would be 2 units away from mean : \[(119-2,~~119+2) ~~~= ~~~(117,~~121)\]